Farmer wants to make a 2400 lb. feed mixture that will be 14% protein.
25% will be Barley(11.7% protein).
the rest or 75 % will consist of oats(11.8% protein) and Soy meal(44.5% protein).
how much of each will he add?
the 2400 lbs. will provide .14 x 2400 = 336 lbs. of prtoein.
the 25% Barley = 600 lbs. it will provide .117 x 600 lbs= 70.20 pounds of protein.
the rest is 1800 lbs. this amount will provide @ 265.80 lbs.
this is where I am stuck.
I don't know how to set up the equation to get the amounts of each of the remaining feeds needed to provide the 265.80 lbs.
Help please.

Answer by askhari139(3)

At this point, you need to divide 1800 lb into oats and soy meal. So, let us assume that farmer adds x lb of oats. Then, the added soy meal will be 1800-x lb.
Total protein content obtained will be (x*0.118)+ ((1800-x)*0.445) which needs to be equal to the 265.80 lb that you calculated.
Therefore,
(x*0.118)+((1800-x)*0.445) = 265.80
1800* 0.445 - 265.80 = x*(0.445-0.118)
535.2 = x*0.327
Hence x = amount of oats added = 1636.69 lb
amount of soy meal = 1800-x = 163.20
CHECK:
total protein required = 0.14*2400 = 336
total protein added = 600*0.117 + 1636.69*0.118 + 163.20*0.445 = 70.2+193.13+72.624 = 335.954